OpenStudy (anonymous):

find the equation of the normal to the curve y=3√(x+1) at the point where x=

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

Normal : The normal at the point P of a curve C is the line through P perpendicular to the tangent at P. This curve is a hyperbola centered at the origin. Solving for y: y = SQRT(5 + x^2) in the first quadrant (since (2, 3) is located in quad I). The derivative is y' = 1/2(5 + x^2)^(-1/2)(2x) The slope of the tangent at (2, 3) is... y'(2) = 1/2(5 + 2^2)^(-1/2)(2)(2) =(1/2)(9)^(-1/2)(4) =2/(9^1/2) = 2/3 The slope of the normal is perpendicular to the tangent, so the slope of the normal is -3/2 Using the point (2, 3), the equation for the normal is (y - 3) / (x - 2) = -3/2

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

where did u get 5+x^2

OpenStudy (shaik0124):

similarly u can solve by placing 3sqrootx+1 instead of 5+x^2

OpenStudy (anonymous):

YEAH

OpenStudy (shaik0124):

method to solve has sshown up by abayomi then u can proceed further